# Fan–Wang type regular black holes in Quasi-Topological Gravity

### v1

@@ I Introduction
Among modified-gravity frameworks, higher-curvature theories are particularly well-motivated because curvature corrections naturally arise in low-energy effective actions of quantum gravity (e.g., string-inspired expansions). While generic higher-derivative theories typically lead to higher-order equations of motion, there exist distinguished classes in which the dynamics remains tractable for highly symmetric ansätze. A prominent example is Lovelock gravity, whose field equations are second order in any dimension, and which contains the Gauss–Bonnet term as the leading nontrivial correction in $D\geq 5$. In five dimensions and higher, another important family is provided by quasi-topological gravity (QTG), in which specific higher-curvature combinations (cubic and beyond) are engineered so that, for static and spherically symmetric spacetimes, the equations reduce to a second-order form while retaining a healthy linearized graviton spectrum around maximally symmetric backgrounds Myers:2010ru . This feature allows one to analyze black hole solutions and their thermodynamics with a level of analytic control comparable to Einstein and Lovelock theories. More recently, these ideas have been extended to broader classes, including generalized quasi-topological gravities and closely related “Einsteinian” higher-curvature theories, which share similar simplifications in symmetric sectors and have been actively studied from both gravitational and holographic viewpoints Bueno:2016xff ; Hennigar:2017ego ; Bueno:2019ycr . It is therefore natural to ask whether such higher-curvature corrections can also accommodate regular black holes, either by themselves or in combination with suitable matter sectors. Indeed, concrete constructions of regular black holes sourced by quasi-topological gravities in $D\geq 5$ have appeared recently Fernandes:2025fnz ; Bueno:2024dgm , providing new examples in which nonsingular interiors arise within a higher-curvature setting. These developments further motivate a systematic investigation of regular black hole solutions in higher dimensions and in gravity theories beyond Einstein’s.

### v2

@@ I Introduction
Among modified-gravity frameworks, higher-curvature theories are particularly well-motivated because curvature corrections naturally arise in low-energy effective actions of quantum gravity (e.g., string-inspired expansions). While generic higher-derivative theories typically lead to higher-order equations of motion, there exist distinguished classes in which the dynamics remains tractable for highly symmetric ansätze. A prominent example is Lovelock gravity, whose field equations are second order in any dimension, and which contains the Gauss–Bonnet term as the leading nontrivial correction in $D\geq 5$. In five dimensions and higher, another important family is provided by quasi-topological gravity (QTG), in which specific higher-curvature combinations (cubic and beyond) are engineered so that, for static and spherically symmetric spacetimes, the equations reduce to a second-order form while retaining a healthy linearized graviton spectrum around maximally symmetric backgrounds Oliva:2010eb ; Myers:2010ru ; Dehghani:2011vu ; Cisterna:2017umf . This feature allows one to analyze black hole solutions and their thermodynamics with a level of analytic control comparable to Einstein and Lovelock theories. More recently, these ideas have been extended to broader classes, including generalized quasi-topological gravities and closely related “Einsteinian” higher-curvature theories, which share similar simplifications in symmetric sectors and have been actively studied from both gravitational and holographic viewpoints Bueno:2016xff ; Hennigar:2017ego ; Bueno:2019ycr ; Aguayo:2025xfi . It is therefore natural to ask whether such higher-curvature corrections can also accommodate regular black holes, either by themselves or in combination with suitable matter sectors. Indeed, concrete constructions of regular black holes sourced by quasi-topological gravities in $D\geq 5$ have appeared recently Fernandes:2025fnz ; Bueno:2024dgm , providing new examples in which nonsingular interiors arise within a higher-curvature setting. This type of regular black hole has recently been constructed in the presence of Maxwell field Hao:2025utc . These developments further motivate a systematic investigation of regular black hole solutions in higher dimensions and in gravity theories beyond Einstein’s.